OpenStudy (anonymous):

Myra's stamp collection consisted of 120 stamps in October. In March, her collection grew to 138 stamps. Find the percent increase between October and March. can someone show me how to work this problem out?

OpenStudy (anonymous):

Let x be the percent. 120 x = 138 x = 138/120 = around 1.85 1.85-1 = 0.85 so around 85% increase

OpenStudy (hoblos):

the increase is 18 stamps 18*100/120 = 15 => the increase percentage is 15%